The perimeter of a rectangle is 242 inches. The length exceeds the width by 55 inches. Find the length and width.
Answer by mananth(16946)

let width be x
length = x+55
Perimeter = 2L+2W
2*x+2(x+55)=242
2x+2x+110=242
4x=132
x=132/4
x=33 , which is the width
Length = x+55
= 33+55=88 inches